OpenStudy (anonymous):

Use elimination to solve the system of equations. Find ordered pairs. { 5x-3y=42 8x+5y=28

OpenStudy (mertsj):

Multiply the top equation by 5 Multiply the second equation by 3 Add the two equations.

OpenStudy (mertsj):

25x-15y=210 24x+15y= 84 49x = 294 Now divide both sides by 49

OpenStudy (mertsj):

What does x = ?

OpenStudy (anonymous):

x=6 right?

OpenStudy (mertsj):

yes. Now find y

OpenStudy (anonymous):

how do i find y?

OpenStudy (mertsj):

Replace x with 6 in any equation that has both an x and a y and solve that equation for y

OpenStudy (anonymous):

so like 5x-3(6)=42?

OpenStudy (mertsj):

I believe you said that x = 6. So why did you replace y with 6 if it is x that is equal to 6?

OpenStudy (anonymous):

oh! so it would be 5(6)-3=42 i don't understand how to do these problems with elimination, I only know how to do them as substitution.

OpenStudy (mertsj):

What happened to the y?

OpenStudy (mertsj):